Gilsan Corporation relevant range of activity is 4,000 units to 8,000 units. When it produces and sells 6,000 units, its average cost per unit is as follows:

                                                                                       AVERAGE COST PER UNIT

DIRECT MATERIALS                            $5.75

DIRECT LABOR                                   $3.00

VARIABLE MANUFACTURING OVERHEAD    $1.60

FIXED MANUFACTURING OVERHEAD              $4.50

FIXED SELLING EXPENSE                              $.075

FIXED ADMINISTRATIVE EXPENSE         $0.60

SALES COMMISSION                                $1.50

VARIABLE ADMINISTRATIVE EXPENSE $0.55

1. FOR FINANCIAL REPORTING PURPOSES, WHAT IS THE TOTAL AMOUNT OF PRODUCT COSTS INCURRED TO MAKE 6,000 UNITS? SHOW ALL WORK

2. FOR FINANCIAL REPORTING PURPOSES WHAT IS THE TOTAL AMOUNT OF PERIOD COSTS TO SELL 6,000 UNITS? SHOW ALL WORK

3. IF 5,000 UNITS ARE SOLD, WHAT IS THE TOTAL AMOUNT OF VARIABLE COSTS RELATED TO THE UNITS SOLD? SHOW ALL WORK

4. IF THE SELLING PRICE IS $19.10 PER UNIT, WHAT IS THE CONTRIBUTION MARGIN PER UNIT SOLD? SHOW ALL WORK

5. WHAT INCREMENTAL MANUFACTURING COST WILL THE COMPANY INCUR IF IT INCREASES PRODUCTION FROM 6,000 TO 6,013 UNITS. SHOW ALL WORK

6. AS A PRODUCT MANAGER OF GLISAN CORP YOU HAVE TO CONSTANTLY MAKE DECISIONS, WHAT METHOD WOULD YOU USE TO REDUCE COST, INCENTIVES EMPLOYEES, AND/OR INCREASE SALES (TRY TO USE ACTUAL NUMBERS TO JUSTIFY YOUR ANSWER)
Answer #1

Product cost = direct material + direct labor + manufacturing overhead

= (5.75+3+1.60+4.50)*6000

=$89100

2.period costs = selling and admin expense

=(0.075+0.60+1.50+0.55)*6000

=$16,350

3.Variable cost =(5.75+3+1.60+1.50+0.55)*5000

=$62,000

4.unit contribution margin = unit selling price- unit variable cost

= 19.10 - 12.4

=$6.7 per unit

5.incremental manufacturing cost = incremental variable manufacturing cost as fixed costs do not increase within the relevant range

=13*(5.75+3+1.60)

=$134.55

6.Contribution margin will be used for these decisions